# Illicit Promotion on Twitter

### v1

@@ 2. Background and Related Works
Different from these previous studies around spamming activities on OSNs, this study focuses on illicit promotions on OSNs (especially Twitter). Compared with general spamming activities, illicit promotion has multiple unique characteristics. Particularly, the goal of illicit promotion is to promote illicit goods/services rather than phishing or scamming the victims. Also, as revealed in this study, miscreants of illicit promotion usually adopt stealthy jargon words and masquerade their promotion content in an evasive manner, which elicits novel technical challenges in terms of detection and understanding.
@@ 3.2. The PIP Analyzer
Category
Description
Pornography
Sexual services, and sexually explicit content, e.g., indecent images and videos.
Gambling
Online or offline gambling services and products.
Illegal Drug
Illegal drugs, e.g., addictive opioid drugs, and prescription drugs.
Surrogacy
Surrogacy services, e.g., surrogate motherhood agencies.
Harassment
Various harassment services, e.g., cyberbullying, stalking, call/sms bombers.
Money Laundering
Various money laundering services, e.g., money muling.
Weapon Sales
The sale of weapons, e.g., P99, a semi-automatic pistol and SR-16, a select-fire rifle.
Data Theft and Leakage
Services offering stolen sensitive datasets, or various hacking tools and services.
Forgery and Fake Documents
Services offering fake or forged documents, e.g., forged passports and fake diplomas.
Crowdturfing
Services of illicit crowdsourcing, e.g., deceptive promotion of the popularity of posts or accounts.
@@ 4. Posts of Illicit Promotion
Hashtags. We observe that benign and popular hashtags have been extensively abused by miscreants to masquerade their PIPs as benign. On the other hand, the main text of many PIPs are benign while the embedded hashtags promote illicit goods or services. Across all categories, PIPs have a median number of 7 hashtags, which is much larger than Twitter’s official recommendation of one or two relevant hashtags (twi, 2023a). More specifically, 73.43% PIPs have embedded three or more hashtags while 49.47% have five or more. Besides, PIP operators may exploit hashtags of trending topics or popular events to enhance their tweets’ reach. For instance, during the FIFA World Cup 2022, illegal gambling operators were found to inject into their PIPs with football hashtags, #WorldCup, #WorldCup2022, or #WorldCupBetting. Similarly, operators of surrogacy PIPs were found to have embedded hashtags closely relevant to the Mid-Autumn Festival, a popular Chinese festival celebrating family reunion and togetherness. Such hashtags include a Chinese hashtag meaning the Mid Autumn and one more meaning family reunion. In addition to increasing visibility, the injection of benign and popular hashtags into PIPs may likely mislead Twitter’s content moderation to some extent. On the other hand, miscreants have composed various malicious hashtags so as to keep the main text of a PIP benign while promoting illicit goods and services in the hashtags, e.g., a Chinese hashtag denoting domestic surrogacy, and a Thai hashtag denoting Cannabis Bangkok, and one more Chinese hashtag denoting mobile phone eavesdropping, etc.
@@ Appendix B The List of Jargon Words as Observed from PIPs
Word
English
Category
Description
蓝精灵
The Smurfs
Drug
Ecstasy
燃料
Fuel
Drug
The alias of marijuana.
叶子
Leaves
Drug
Marijuana
飞叶子
Flying Leaves
Drug
Smoke marijuana
飞行
Fly
Drug
The Behavior of smoking marijuana.
机长
Pilot
Drug
Drug dealer selling marijuana.
农夫
Farmer
Drug
Man who grows marijuana.
鲍鱼
Abalone
Pornography
Female genitalia
遛鸟
Walking the birds
Pornography
Men relaxing their genitals in public or outdoors
呦呦
Yoyo
Pornography
Used to describe child porn.
四件套
Four pieces set
Data Leakage
Sell the ’four-piece set’ of bank cards (mobile phone card, bank card, U-disk, and photocopy of ID card).
跑分
Benchmarking
Money Laundering
Using a third-party payment account to collect funds on behalf of others, and then transferring the funds to earn a commission.
渔夫
Fisherman
Money Laundering
Phishing using own personal number, responsible for finding victims
船长
Captain
Money Laundering
Responsible for setting up phishing sites.
黄河
Yellow River
Weapon
A brand of gun
海螺
Conch
Gambling
Board and card game and electronic game system.

### v2

@@ 2. Background and Related Works
Different from these studies, this study focuses on illicit promotions on OSNs (especially Twitter). Also, In contrast to OSN spamming, illicit promotion aims to advertise illicit goods/services rather than phishing or scamming the victims.
Illicit promotion on OSNs. Previous works have also made efforts to reveal illicit promotion on OSNs, but most of them focus on a specific category or entity instead of comprehensive detection and understanding. For example, authors of (Kalyanam and Mackey, 2017; Katsuki et al., 2015) paid attention to illicit pharmacies on Twitter and (Yuan et al., 2019) studied adversarial pornography images crawled from popular OSNs along with the underground business behind them. In this paper, we step forward to build up a general tool chain to detect and analyze PIPs along with efficacy demonstrated for diverse PIP categories and multilingual PIP instances. Furthermore, most findings distilled from this study are applicable to the ecosystem of PIPs rather than being limited to a specific PIP category.
@@ 3.2. The PIP Analyzer
Category Description
Pornography Sexual services, and sexually explicit content, e.g., indecent images and videos.
Gambling Online or offline gambling services and products.
Illegal Drug Illegal drugs, e.g., addictive opioid drugs, and prescription drugs.
Surrogacy Surrogacy services, e.g., surrogate motherhood agencies.
Harassment Various harassment services, e.g., cyberbullying, stalking, call/sms bombers.
Money Laundering Various money laundering services, e.g., money muling.
Weapon Sales The sale of weapons, e.g., P99, a semi-automatic pistol and SR-16, a select-fire rifle.
Data Theft and Leakage Services offering stolen sensitive datasets, or various hacking tools and services.
Forgery and Fake Documents Services offering fake or forged documents, e.g., forged passports and fake diplomas.
Crowdturfing Services of illicit crowdsourcing, e.g., deceptive promotion of the popularity of posts or accounts.
@@ 4. Posts of Illicit Promotion
PIP engagements. Over time, PIPs accumulate engagements such as likes, replies, retweets and quotes (i.e. retweets with comments). Given the date a PIP is published $t_{p}$ and the date it is crawled $t_{c}$, $t_{e}=t_{c}-t_{p}$ is the elapse time of a PIP, i.e., the days passed since it is posted. As described in §5, 90% PIPs survive the first two months, thus we group PIPs based on elapse time ranging from 1 day to 60 days, and calculate the average engagement of each group. As shown in Figure 6, PIPs indeed receive a non-negligible volume of engagement, e.g., PIPs can receive 374 likes on average after being posted for 30 days, while the average number of likes a regular tweet can receive is 37 (ave, 2021). What is also observed is that the extent of engagement declines gradually after 35 days. We believe it is because PIPs with more engagement are more likely to be detected and unpublished by Twitter, which renders the survivorship bias wherein only the surviving cases will be counted for groups of long elapse time. However we are not clear about the reason why average retweets show significant fluctuations after 40 days.
Figure 6. Average engagements of PIPs over time.
@@
Hashtags. On one hand, benign and popular hashtags are extensively abused by miscreants to masquerade their PIPs. Across all tweets we’ve collected, non-PIPs have a median number of 0 hashtags and an average number of 2.27 hashtags, while it is 5 and 6.98 for PIPs respectively, which is much larger than Twitter’s official recommendation of one or two relevant hashtags (twi, 2023a). Specifically, 73.43% PIPs have embedded three or more hashtags while 49.47% have five or more. Besides, PIP operators may exploit hashtags of trending topics or popular events to enhance their tweets’ reach. For instance, during the FIFA World Cup 2022, illegal gambling operators were found to inject into their PIPs with football hashtags, #WorldCup, #WorldCup2022, or #WorldCupBetting. Similarly, operators of surrogacy PIPs were found to have embedded hashtags closely relevant to the Mid-Autumn Festival, a popular Chinese festival celebrating family reunion and togetherness. Such hashtags include a Chinese hashtag meaning the Mid Autumn and one more meaning family reunion. In addition to increasing visibility, the injection of benign and popular hashtags into PIPs may likely mislead Twitter’s content moderation to some extent. On the other hand, miscreants compose various malicious hashtags so as to keep the main text of a PIP benign while promoting illicit goods and services in the hashtags, e.g., a Chinese hashtag denoting domestic surrogacy, and a Thai hashtag denoting Cannabis Bangkok, and one more Chinese hashtag denoting mobile phone eavesdropping.
@@ Appendix B The List of Jargon Words as Observed from PIPs
Word English Category Description
蓝精灵 The Smurfs Drug Ecstasy
燃料 Fuel Drug The alias of marijuana.
叶子 Leaves Drug Marijuana
飞叶子 Flying Leaves Drug Smoke marijuana
飞行 Fly Drug The Behavior of smoking marijuana.
机长 Pilot Drug Drug dealer selling marijuana.
农夫 Farmer Drug Man who grows marijuana.
鲍鱼 Abalone Pornography Female genitalia
遛鸟 Walking the birds Pornography Men relaxing their genitals in public or outdoors
呦呦 Yoyo Pornography Used to describe child porn.
四件套 Four pieces set Data Leakage Sell the ’four-piece set’ of bank cards (mobile phone card, bank card, U-disk, and photocopy of ID card).
跑分 Benchmarking Money Laundering Using a third-party payment account to collect funds on behalf of others, and then transferring the funds to earn a commission.
渔夫 Fisherman Money Laundering Phishing using own personal number, responsible for finding victims
船长 Captain Money Laundering Responsible for setting up phishing sites.
黄河 Yellow River Weapon A brand of gun
海螺 Conch Gambling Board and card game and electronic game system.